# Chow and augmented Chow polynomials as evaluations of Poincaré-extended $\mathbf{a}\mathbf{b}$-indices

### v2

@@ 2.3. Evaluating the Poincaré-extended $\mathbf{a}\mathbf{b}$-index
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\textsf{H}_{P}(x)\,,$
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\underline{\textsf{H}}_{P}(x)\,.$
@@
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)=\sum_{\mathcal{F}}\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$
where the sum ranges over all maximal chains $\mathcal{F}$ in $P$. Applying the evaluation $y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x$ to the substitution $\omega$ gives that every occurrence in ${\sf u}_{\mathcal{F},\emptyset}$ of the pattern $\mathbf{a}\mathbf{b}$ becomes a factor $x-x^{2}-x^{2}+x^{3}=x(1-x)^{2}$, every remaining $\mathbf{a}$ yields a factor $1-x^{2}=(1+x)(1-x)$, and every remaining $\mathbf{b}$ yields a factor 00. We therefore obtain
@@
if $\lambda_{\mathcal{F}}$ contains a descent that does not immediately follow an ascent, and otherwise
$\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{|}_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$ $=(1-x)^{n}\cdot x^{\mathsf{des}(\lambda_{\mathcal{F}})}(1+x)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
Summing over all maximal chains $\mathcal{F}$ in $P$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$ then gives
${\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
@@
The argument to deduce Theorem 1.1 is almost the same. The only difference is that now every initial variable is removed from every $\mathbf{a}\mathbf{b}$-monomial to obtain ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ from ${\text{\scriptsize\sf ex}}\Psi_{P}(x,\mathbf{a},\mathbf{b})$. This is, we have
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big{(}{\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big{)}=\sum_{\mathcal{F}}\iota\big{(}\omega\big{(}{\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big{)}\big{)}\,.$
@@
Summing now over all maximal chains $\mathcal{F}$ in $P$ such that for $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ it holds that $\lambda_{1}<\lambda_{2}$ and also that $\lambda_{i}>\lambda_{i+1}$ implies $\lambda_{i-1}<\lambda_{i}$ for $i\in\{2,\dots,n-1\}$ then gives
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\lambda_{\mathcal{F}})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\lambda_{\mathcal{F}})}\,.$
@@
$=x^{n}(x-1)^{-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x^{-1},\tfrac{x}{x-1},\tfrac{1}{x-1})$
where we used in the last equation that ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ has exactly $n-1$ factors, either being $\mathbf{b}$ for ranks in $\mathcal{C}\setminus\{\mathcal{C}_{k+1}\}$ or being $\mathbf{a}-\mathbf{b}$ for ranks not in $\mathcal{C}$, and moreover that evaluating $\mathbf{b}\mapsto\frac{1}{x-1}$ and $\mathbf{a}\mapsto\frac{x}{x-1}$ evaluates $\mathbf{a}-\mathbf{b}\mapsto 1$ and at the same time yields a factor $(x-1)^{1-k_{\mathcal{C}}}$. Using the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$, we obtain
$\underline{\textsf{H}}_{P}(x)$ $=x^{-1}(x^{-1}-1)^{-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})$
$=(1-x)^{-n}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)$
where we used that ${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ is homogeneous of $\mathbf{a}\mathbf{b}$-degree $n-1$ and thus
${\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})=(1-x)^{n-1}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(-x,1,x)\,.$
@@
$\textsf{H}_{P}(x)=(1-x)^{-n}{\text{\scriptsize\sf ex}}\Psi_{P}(-x,1,x)$
@@
We used in the proof the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$. As the symmetry of the evaluated extended $\mathbf{a}\mathbf{b}$-index is clear from its combinatorial description, this is not a prerequired property of the Chow polynomial and we only used it here for the clarity of the argument.
@@ 2.4. Evaluating the coarse flag Hilbert-Poincaré series
The coarse flag Hilbert-Poincaré series ${\sf cfHP}_{P}(y,t)$ is originally defined and studied in [MV23] as an algebraic invariant in geometric group theory. It is shown in [DMS23, Corollary 2.22] that this rational function can be expressed as
${\sf cfHP}_{P}(y,t)=(1-t)^{-n}{\text{\scriptsize\sf ex}}\widetilde{\Psi}_{P}(y,1,t)\,.$
@@
$\underline{\textsf{H}}_{P}(x)={\sf cfHP}_{P}(x,-x)\,.$
Applying $\leavevmode\nobreak\ \iota$ to ${\text{\scriptsize\sf ex}}\Psi_{P}(y,\mathbf{a},\mathbf{b})$ restricts the situation to chains starting in $\hat{0}$ as in Proposition 2.3. It is thus not immediately possible to also obtain the augmented Chow polynomial from the coarse flag Hilbert-Poincaré series.
3. Chow and augmented Chow polynomials for braid arrangement
In this section, we apply the main theorems to obtain a compact expansion of the Chow and the augmented Chow polynomials for the braid arrangement $\mathcal{B}_{n}$ given by the hyperplanes
[…]
in $V=\big{\{}(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\mid x_{1}+\dots+x_{n+1}=0\big{\}}\cong\mathbb{R}^{n}$.
@@ 3. Chow and augmented Chow polynomials for braid arrangement
where the sum ranges over all tuples $(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that $a_{1}\leq a_{2}$ and $a_{i}>a_{i+1}$ implies $a_{i-1}\leq a_{i}$ for $i\in\{2,\dots,n-1\}$. We moreover have
@@
where the sum ranges over all tuples $(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that $a_{i}>a_{i+1}$ implies $a_{i-1}\leq a_{i}$ for $i\in\{2,\dots,n-1\}$.
@@
Let now $n=3$. For the Chow polynomials, the sum ranges over $\{(1,1,1),(1,2,1),(2,2,1)\}$ and we obtain
@@
[…]
@@
This $R$-labeling was recently used in [HS24] to derive an explicit form for the coarse flag Hilbert-Poincaré series of the braid arrangement. In light of Corollary 2.8, it seems natural that the max-to-min labeling also plays a crucial role in the understanding of the Chow polynomials of the braid arrangements.
To deduce Theorem 3.1 from Theorems 1.1, LABEL: and 1.2, we calculate for a given permutation $\sigma$ of $\{2,\dots,n+1\}$ the number of maximal chains $\mathcal{L}$ in $\Pi_{n}$ for which $\lambda_{\mathcal{F}}=\sigma$. (Our taylormade version of) the inversion sequence $\mathsf{Inv}(\sigma)=(a_{1},\dots,a_{n})$ of a permutation $\sigma=[\sigma_{1},\dots,\sigma_{n}]$ of $\{2,\dots,n+1\}$ in one-line notation is given by
@@
Observe that the entry $a_{i}$ in the inversion sequence $\mathsf{Inv}(\sigma)$ may be computed from the partial permutation $[\sigma_{1},\dots,\sigma_{i}]$ as
@@
#$\#\{\text{maximal chains }\mathcal{F}\text{ in }\Pi_{n}\mid\lambda_{\mathcal{F}}=\sigma\}=a_{1}\cdot\ldots\cdot a_{n}\,.$
@@
Conversely, the labels after reaching $\mathbf{B}$ are $\{\min B_{1},\dots,\min B_{\ell}\}\setminus\{1\}$. Moreover, the labeled interval between $\mathbf{B}$ and the maximal element $\{1,\dots,n+1\}\in\Pi_{n}$ does not depend on $\mathbf{B}$ itself as it is simply the set partition lattice on the set $\{\min B_{1},\dots,\min B_{k}\}$ again with the max-of-min labeling.
Let now $i\in\{1,\dots,n\}$, let $\sigma=[\sigma_{1},\dots,\sigma_{i}]$ be a non-empty partial permutation, i.e., a list of $i$ distinct numbers in $\{2,\dots,n+1\}$ and let $\sigma^{\prime}=[\sigma_{1},\dots,\sigma_{i-1}]$ be the partial permutation obtained by removing from $\sigma$ the last entry. Set $a_{i}$ to be the $i$’s entry in the inversion sequence of any permutation starting with $\sigma$ as given in (5). Finally, denote by $m_{\sigma}$ the number of maximal chains $\mathcal{F}$ in $\Pi_{n}$ for which the labeling $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ starts with $(\lambda_{1},\dots,\lambda_{i})=(\sigma_{1},\dots,\sigma_{i})$ and by $m_{\sigma^{\prime}}$ the number of maximal chains for which the labeling starts with $(\sigma_{1},\dots,\sigma_{i-1})$. Then
$m_{\sigma}=m_{\sigma^{\prime}}\cdot a_{i}\,,$

### v3

@@ 2.3. Evaluating the Poincaré-extended $\mathbf{a}\mathbf{b}$-index
${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\textsf{H}_{P}(x)\,$,
${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,1,x)$ $=(1-x)^{n}\cdot\underline{\textsf{H}}_{P}(x)\,.$
@@
${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(-x,1,x)=\sum_{\mathcal{F}}\omega\big({\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big)\big|_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$
[…]
@@
if ${\sf Des}(\mathcal{F})$ is not isolated, and otherwise
$\omega\big({\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big)\big|_{y=-x,\ \mathbf{a}=1,\ \mathbf{b}=x}$ $=(1-x)^{n}\cdot x^{\mathsf{des}(\mathcal{F})}(1+x)^{n-2\ \!\!\mathsf{des}(\mathcal{F})}\,.$
Summing over all maximal chains $\mathcal{F}$ in $P$ such that for ${\sf Des}(\mathcal{F})$ is isolated then gives
(5) ${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\mathcal{F})}(x+1)^{n-2\ \!\!\mathsf{des}(\mathcal{F})}\,.$
@@
The argument to deduce Theorem 1.1 is almost the same. The only difference is that now every initial variable is removed from every $\mathbf{a}\mathbf{b}$-monomial to obtain ${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ from ${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})$. This is, we have
${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})=\iota\big({\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})\big)=\sum_{\mathcal{F}}\iota\big(\omega\big({\sf u}_{\mathcal{F},\emptyset}(\mathbf{a},\mathbf{b})\big)\big)\,.$
@@
Summing now over all maximal chains $\mathcal{F}$ in $P$ such that ${\sf Des}(\mathcal{F})$ is isolated and $1\notin{\sf Des}(\mathcal{F})$ then gives
${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,1,x)=(1-x)^{n}\sum_{\mathcal{F}}x^{\mathsf{des}(\mathcal{F})}(x+1)^{n-1-2\ \!\!\mathsf{des}(\mathcal{F})}\,.$
@@
$=x^{n}(x-1)^{-1}{\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x^{-1},\tfrac{x}{x-1},\tfrac{1}{x-1})$
where we used in the last equation that ${\sf wt}_{\mathcal{C}}(\mathbf{a},\mathbf{b})$ has exactly n1n-1 factors, either being $\mathbf{b}$ for ranks in $\mathcal{C}\setminus\{\mathcal{C}_{k+1}\}$ or being $\mathbf{a}-\mathbf{b}$ for ranks not in $\mathcal{C}$, and moreover that evaluating $\mathbf{b}\mapsto\frac{1}{x-1}$ and $\mathbf{a}\mapsto\frac{x}{x-1}$ evaluates $\mathbf{a}-\mathbf{b}\mapsto 1$ and at the same time yields a factor $(x-1)^{1-k_{\mathcal{C}}}$. Using the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$, we obtain
$\underline{\textsf{H}}_{P}(x)$ $=x^{-1}(x^{-1}-1)^{-1}{\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})$
(6) $=(1-x)^{-n}{\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,1,x)$
where we used that ${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(y,\mathbf{a},\mathbf{b})$ is homogeneous of $\mathbf{a}\mathbf{b}$-degree n1n-1 and thus
${\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,\tfrac{1}{1-x},\tfrac{x}{1-x})=(1-x)^{-(n-1)}\cdot{\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(-x,1,x)\,.$
@@
$\textsf{H}_{P}(x)=(1-x)^{-n}{\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(-x,1,x)$
@@
In the proof, we used the symmetry $\underline{\textsf{H}}_{P}(x)=x^{n-1}\underline{\textsf{H}}_{P}(x^{-1})$. As the symmetry of $(1-x)^{-n}{\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(-x,1,x)$ is clear from its combinatorial description established in (5), this is not a required property of the Chow polynomial as we could instead have used the symmetry of the right-hand side of (6). We also observe that we have not made use of the property that the poset admits an $R$-labeling. This evaluation thus holds for all finite graded posets.
@@ 2.4. Evaluating the coarse flag Hilbert-Poincaré series
The coarse flag Hilbert-Poincaré series ${\sf cfHP}_{P}(y,t)$ is originally defined and studied in [MV24] as an algebraic invariant in geometric group theory. It is shown in [DBMS25, Corollary 2.22] that this rational function can be expressed as
${\sf cfHP}_{P}(y,t)=(1-t)^{-n}{\text{\scriptsize\sf ex}\hskip-1.0pt}\widetilde{\Psi}_{P}(y,1,t)\,.$
@@
$\underline{\textsf{H}}_{P}(x)={\sf cfHP}_{P}(-x,x)\,.$
Applying $\penalty 10000\ \iota$ to ${\text{\scriptsize\sf ex}\hskip-1.0pt}\Psi_{P}(y,\mathbf{a},\mathbf{b})$ restricts the situation to chains starting in $\hat{0}$ as in Proposition 2.3. It is thus not immediately possible to also obtain the augmented Chow polynomial from the coarse flag Hilbert-Poincaré series.
3. Chow and augmented Chow polynomials
for braid arrangement
In this section, we apply the main theorems to obtain explicit closed formulas of the Chow and the augmented Chow polynomials for the braid arrangement $\mathcal{B}_{n}$ given by the hyperplanes
$H_{ij}=\left\{x_{i}=x_{j}\;|\;1\leq i<j\leq n+1\right\}$
in $V=\big\{(x_{1},\dots,x_{n+1})\in\mathbb{R}^{n+1}\mid x_{1}+\dots+x_{n+1}=0\big\}\cong\mathbb{R}^{n}$. Inductive formulas for all reflection arrangements of classical types have been established in [GS12].
@@ 3. Chow and augmented Chow polynomials
where the sum ranges over all $a=(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that ${\sf Des}(a)$ is isolated and $1\notin{\sf Des}(a)$. We moreover have
@@
where the sum ranges over all $a=(a_{1},\dots,a_{n})$ with $a_{i}\in\{1,\dots,n+1-i\}$ such that ${\sf Des}(a)$ is isolated.
@@
Let now $n=3$. For the Chow polynomials, the sum ranges over $\{(1,1,1)$, $(1,2,1)$, (2,2,1)}(2,2,1)\} and we obtain
@@
and for the augmented Chow polynomials, the sum ranges over $\{(1,1,1)$, $(1,2,1)$, $(2,2,1)$, $(2,1,1)$, (3,1,1)}(3,1,1)\} and we obtain
@@
These calculations agree with the results presented in [GS12, Section 3.2].
@@
This $R$-labeling was recently used in [HS24] to derive an explicit form for the coarse flag Hilbert-Poincaré series of the braid arrangement. In light of Corollary 2.8, it seems natural that the max-of-min labeling also plays a crucial role in the understanding of the Chow polynomials of the braid arrangements. Moreover, Brändén and Vecchi recently proved the real-rootedness of the augmented Chow polynomial of the braid arrangement in [BV25].
To deduce Theorem 3.1 from Theorems 1.1, LABEL: and 1.2, we calculate for a given permutation $\sigma$ of $\{2,\dots,n+1\}$ the number of maximal chains $\mathcal{L}$ in $\Pi_{n}$ for which $\lambda_{\mathcal{F}}=\sigma$. (Our tailor-made version of) the inversion sequence $\mathsf{Inv}(\sigma)=(a_{1},\dots,a_{n})$ of a permutation $\sigma=[\sigma_{1},\dots,\sigma_{n}]$ of $\{2,\dots,n+1\}$ in one-line notation is given by
@@
Observe that the entry $a_{i}$ in the inversion sequence $\mathsf{Inv}(\sigma)$ may be computed from the partial permutation $\sigma_{1},\dots,\sigma_{i}]$ as
@@
$\{\text{maximal chains }\mathcal{F}\text{ in }\Pi_{n}\mid\lambda_{\mathcal{F}}=\sigma\}=a_{1}\cdot\ldots\cdot a_{n}\,.$
@@
Conversely, the labels after reaching $\mathbf{B}$ are $\{\min B_{1},\dots,\min B_{\ell}\}\setminus\{1\}$. Moreover, the labeled interval between $\mathbf{B}$ and the maximal element $\{1,\dots,n+1\}\in\Pi_{n}$ does not depend on $\mathbf{B}$ itself as it is simply the set partition lattice on the set $\{\min B_{1},\dots,\min B_{\ell}\}$ again with the max-of-min labeling.
Let now $i\in\{1,\dots,n\}$, let $\sigma=[\sigma_{1},\dots,\sigma_{i}]$ be a non-empty partial permutation, i.e., a list of $i$ distinct numbers in $\{2,\dots,n+1\}$ and let $\sigma^{\prime}=[\sigma_{1},\dots,\sigma_{i-1}]$ be the partial permutation obtained by removing from $\sigma$ the last entry. Set $a_{i}$ to be the $i$’s entry in the inversion sequence of any permutation starting with $\sigma$ as given in (7). Finally, denote by $m_{\sigma}$ the number of maximal chains $\mathcal{F}$ in $\Pi_{n}$ for which the labeling $\lambda_{\mathcal{F}}=(\lambda_{1},\dots,\lambda_{n})$ starts with $(\lambda_{1},\dots,\lambda_{i})=(\sigma_{1},\dots,\sigma_{i})$, meaning that $m_{\sigma}$ counts initial chains $\mathcal{C}=(\mathcal{C}_{0},\dots,\mathcal{C}_{i})$ such that ${\sf rank}(\mathcal{C}_{j})=j$ and $\lambda_{j}=\sigma_{j}$ for all $j\in\{1,\dots,i\}$. Similarly, denote by $m_{\sigma^{\prime}}$ the number of maximal chains for which the labeling starts with $(\sigma_{1},\dots,\sigma_{i-1})$. Then
$m_{\sigma^{\prime}}=m_{\sigma}\cdot a_{i}\,$,